Please help me with these answers. Thanks

7. Listed below are the average prices of MyFavoriteStock, Inc. since 2000. Use your calculator to find (a) the Regression Formula (assume it is a linear function: LinReg) and (b) the predicted value in 2020.
Years since 2000 1 2 3 4 5 6
Stock Value per share $8 $10 $13 $11 $14 $16
Regression Formula = __________________
2020 predicted price = __________________

Answer by uma(370)

The Regression equation of Y on X is given by Y = a + bX
Considering the years 1,2,3,4,5,6 as X and the prices 8,10,13,11,14,16 as Y.
we have sum(Y) = Na + b*sum(x)
and sum(XY) = a*sum(X) + b*sum(X^2)
From the given values , sum(X) = 21, sum(Y) = 72, sum(XY)= 277, sum(X^2) = 91
and N = 6
So 72 = 6a + 21b and
24 = 2a + 7b
Solving the above two equations we get a = 7 and b = 10/7 = 1.4
Thus the regression formula is Y = 7 + 1.4X
Now when X = 20, Y = 7 + 1.4 * 20
Y = 7 + 28
= 35
So 2020 predicted price would be $35

Good Luck!!!